### v1

@@ 2.1 Food Recommendation
In response to the diversified needs for food, FRSs are committed to providing accurate and personalized recommendation services for users. Recently, leveraging recipe metadata, such as ingredients and images, within the general CF framework has proven effective in mitigating data sparsity challenges and refining user preference modeling. HAFR [1] develops a hierarchical attention mechanism to incorporate the feature interaction between metadata for recommendation. Moving beyond traditional methods, FGCN [2] explores leveraging graph neural networks to capture high-order relationships through graph propagation and aggregation. Furthermore, SCHGN [13] captures the user’s preference on food calories with self-supervised heterogeneous graph network. In addition, RecipeRec [15] constructs heterogeneous recipe graphs to model the structure. GreenRec [18] accommodates health criteria into recommender to promote sustainable lifestyles.
@@ 2.2 Multi-modal Recommendation
Multi-modal recommendation systems (MMRSs) strive to enhance items’ representation for better recommendation performance by incorporating features from different modalities, including textual content and visual content [25, 19, 20, 21, 22]. Early studies, such as VBPR [5], incorporate multi-modal contents with items’ ID embeddings to extend the general CF framework. Inspired by the great success of graph-based recommendation methods [6, 31, 28, 17, 24], MMGCN [16] attempts to inject high-order semantics into user/item representations via message passing on modality-specific graphs. From another perspective, FREEDOM [30] constructs item-item semantic graph from pre-trained multi-modal features to supplement the user-item interactions. An emerging trend is the integration of self-supervised learning methods into MMRSs, showcasing a notable boost. BM3 [32] and LGMRec [3] improve user and item representation learning by adopting contrastive learning objective. DRAGON [26] utilizes homogeneous graphs based on item multi-modal information and a heterogeneous bipartite user-item graph for effective recommendation. DGVAE [29] disentangles the user-item interactions and multi-modal information with graph learning.
@@ 4.3 Cross-modal Self-supervised Learning
where $dCov(\cdot)$ is the distance covariance between two matrices, and $dVar(\cdot)$ represents its own distance covariance.

### v2

@@ 2.1 Food Recommendation
In response to the diversified needs for food, FRSs are committed to providing accurate and personalized recommendation services for users. Recently, leveraging recipe metadata, such as ingredients and images, within the general CF framework has proven effective in mitigating data sparsity challenges and refining user preference modeling. HAFR [1] develops a hierarchical attention mechanism to incorporate the feature interaction between metadata for recommendation. Moving beyond traditional methods, FGCN [2] explores leveraging graph neural networks to capture high-order relationships through graph propagation and aggregation. Furthermore, SCHGN [13] captures the user’s preference on food calories with self-supervised heterogeneous graph network. In addition, RecipeRec [15] constructs heterogeneous recipe graphs to model the structure. Some recent works  [25, 18] accommodates health criteria into food recommender to promote sustainable lifestyles [34].
@@ 2.2 Multi-modal Recommendation
Multi-modal recommendation systems (MMRSs) strive to enhance items’ representation for better recommendation performance by incorporating features from different modalities, including textual content and visual content [27, 19, 20, 21, 22]. Early studies, such as VBPR [5], incorporate multi-modal contents with items’ ID embeddings to extend the general CF framework. Inspired by the great success of graph-based recommendation methods [6, 33, 30, 17, 24, 26], MMGCN [16] attempts to inject high-order semantics into user/item representations via message passing on modality-specific graphs. From another perspective, FREEDOM [32] constructs item-item semantic graph from pre-trained multi-modal features to supplement the user-item interactions. An emerging trend is the integration of self-supervised learning methods into MMRSs, showcasing a notable boost. BM3 [35] and LGMRec [3] improve user and item representation learning by adopting contrastive learning objective. DRAGON [28] utilizes homogeneous graphs based on item multi-modal information and a heterogeneous bipartite user-item graph for effective recommendation. DGVAE [31] disentangles the user-item interactions and multi-modal information with graph learning.
@@ 4.3 Cross-modal Self-supervised Learning
where $dCov(\cdot)$ is the distance covariance between two matrices, and $dVar(\cdot)$ represents its own distance variance. Distance correlation helps mitigate the representation collapse often induced by other self-supervised constraints by maximizing dependencies between different modality-specific graphs.
@@ 6 CONCLUSION
6.0.1 Acknowledgements.
This research is supported, in part, by the Natural Science Foundation of China (No.92367202, No.62202279), Outstanding Youth Science Fund Project (Overseas) of Shandong (No.2023HWYQ-039), Natural Science Foundation of Shandong Province (No.ZR2022QF018, No.ZR2023LZH006), the Youth Student Fundamental Research Funds of Shandong University, the State Scholarship Fund by the China Scholarship Council (CSC). This research is also supported, in part, by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).